OpenStudy (anonymous):

5n^2 + 19n +12 How do I factor this trinomial step by step?

OpenStudy (campbell_st):

well there are a lot of different methods I use this as a fail-safe method for a quadratic \[ax^2 + bx + c \] multiply the coefficient of the leading term by the constant... a x c 5 x 12 = 60 find the factors of 60 that add to 19 15, 4 I then have the \[\frac{(ax + factor1)(ax + fractor2)}{a}\] in your question its \[\frac{(5x + 15)(5x + 4)}{5} \] (5x + 15) can be written as 5(x + 3) so you have \[\frac{5(x + 3)(5x + 4)}{5}\] cancel the common factor of 5 from the numerator and denominator which gives \[5x^2 + 19x + 12 = (x + 3)(5x + 4)\] hope this helps